OpenStudy (anonymous):

for the following chemical reaction, what mass of calcium carbonate ( in grams) will be produced from 2.41 mol of sodium carbonateNa2Co3+Ca(NO3)2→CaCO3+2NaNO3

OpenStudy (aaronq):

The reactant is already in moles.. so now, make a ratio of the two compounds involved. Set it up so that the moles for each are over their respective stoichiometric coefficients \(\sf \dfrac{moles~of~sodium ~carbonate }{sodium ~carbonate's~coefficient}=\dfrac{moles~of~calcium ~carbonate}{calcium~carbonate's~coefficient}\)

OpenStudy (aaronq):

After, solve algebraically. Next use the moles and the molar mass to find the mass. \(\sf moles=\dfrac{mass}{Molar~mass}\)